# How LLMs Learn to Reason: A Complex Network Perspective

### v1

@@ 3 The Double-Edged Sword of Sparsity: from catastrophic Forgetting to Simulated Annealing scheme
From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability. It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with (Li et al., 2024). Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I). As expected, the intervention triggered a precipitous collapse in overall performance in both CoNet and the 1.5B LLM (Fig. 44(a), 4(b)). Having reproduced this known macroscopic phenomenon, we leveraged CoNet’s transparency to inspect the microscopic aftermath.
The result is a direct validation of our structural explanation. The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web. Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, surgically severing the single connection points for vast downstream subgraphs. Fig. 44(c)4(d) provides a stark visualization of this: a unified web is instantly fragmented, offering a clear, structural cause for the catastrophic performance drop.
@@ 4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
Our sparse web theory reveals that the smooth, aggregate reward curve of RLVR training masks a dramatic microscopic behavior. Individual problem trajectories in Fig. 5 reveal a fundamental tension born from the network’s sparse topology: a subtractive process of frustration-induced forgetting, peaking at the onset of slow learning, and an opposing, additive process of phase-transition-like learning that drives knowledge integration throughout the slow-learning stage.
@@
Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality. (i) Frustration-Induced Forgetting: At the transition to slow learning (see inset in (a)), the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed. (ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green). The smoother gradients in the LLM (vs. the clean steps in CoNet) can be attributed to a classic finite-size effect, as explained in the main text.
@@ 4.1 Frustration-Induced Forgetting: The Price of Integration
The first dynamic is driven by frustration-induced forgetting: a competitive process where the advancement of some skills forces the deterministic regression of others as the model integrates its “skill islands”. This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig. 5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025).
This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where skill islands compete for limited connections under the $\left<k\right>\approx 2$ constraint. Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig. 6(b), blue). This subsequent degradation in performance is consistent with the phenomenon of policy collapse, which has been recently reported in several studies to occur after extended RLVR, leading to a loss of exploratory power (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
@@ 4.2 Phase-Transition-Like Learning: The Engine of Growth
The opposing dynamic, phase-transition-like learning, is an important engine of new knowledge acquisition (Power et al., 2022; Lin et al., 2025). The evidence for this is clearly shown in CoNet’s sharp, step-like accuracy jumps (Fig. 5(a)), and the steep accelerated learning gradients observed in the 1.5B LLM (Fig. 5(b), more details in Appendix G). This phenomenon is a multi-task extension of the "Learning at Criticality" (LaC) mechanism, first identified in a single-task context (Cai et al., 2025). The LaC framework models this process as a continuous phase transition(Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one. Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions. This context explains the smoother gradient in the LLM. Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of a finite-size effect(Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting. Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition. Shielded from interference, a new skill can be efficiently “bolted on” via a clean phase transition. This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025). We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
5 Annealed-RLVR: An Algorithm Born from the Sparse Web Theory
(a) Randomly Selected 512 Training Problems
@@ 5 Annealed-RLVR: An Algorithm Born from the Sparse Web Theory
Figure 6: Annealed-RLVR Outperforms Standard RLVR. This figure compares our Annealed-RLVR (“Annealed”) with a standard RLVR baseline (“RLVR”) on (a) in-distribution problems and (b) the out-of-distribution Minerva dataset. (Right panels) The best@k accuracy curves show that Annealed-RLVR achieves superior performance in both settings, confirming improved generalization. (Left panels) The histograms reveal the underlying mechanism: the annealing intervention systematically reduces the population of unsolved problems (low-accuracy bins) and improves the mastered ones (high-accuracy bins).
@@
The algorithm operationalizes this concept with a precisely timed intervention, triggered by macroscopic signatures of the maximally frustrated state. At this juncture, a brief SFT “heating” phase begins. The timing is critical: because the “skill islands” have not yet solidified into a single web, the system is uniquely robust to a targeted shock that would otherwise cause catastrophic forgetting—a theoretical claim confirmed by our experiments (Fig. 4). The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes. We implement this by surgically applying SFT only to problems with low accuracy (e.g., < 0.1) for which correct solutions exist. Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
The results validate this theory-driven approach (Fig. 6). Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution problems and, crucially, the out-of-distribution Minerva dataset, confirming the intervention fosters true generalization. The histograms reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration. The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery. This results in a systematic reduction of unsolved problems and a corresponding population shift towards solved ones, confirming that our intervention guides the model past the inherent bottlenecks of standard RLVR to achieve superior reasoning. More details are given in Appendix H and I.
@@ 6 Conclusion and Outlook
A natural consequence of this sparse structure is a fundamental RLVR mechanism for forgetting and learning within the microscopic training process. We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently “bolted on” at the web’s expanding frontier.
@@ 7 Related Work
8 ACKNOWLEDGMENTS
The authors thank Haijun Zhou, Lei Wang, Hong Zhao and Ganqu Cui for inspiring discussions. K. C. and X. C. are supported by the National Key Research and Development Program of China under Grant No. 2024YFA1408604 and the National Natural Science Foundation of China under Grants No. 12047503 and No. 12447103. P. Z. is supported by Project 12325501 of the National Natural Science Foundation of China. Y. D. and S. H. are supported by the National Natural Science Foundation of China under Grants No. 12275263, the Innovation Program for Quantum Science and Technology under Grant No. 2021ZD0301900, and the Natural Science Foundation of Fujian Province of China under Grant No. 2023J02032. Z. Y. is supported by the National Natural Science Foundation of China under Grants No. 12304288 and No. 12247101.
@@ Appendix B Extended Related Work
System-2 Reasoning in LLMs. The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior. While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting more robust, System-2 deliberation. Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps. Other works aim to build this deliberative capacity into the model’s architecture or training process through self-correction and reflection (Shinn et al., 2023; Madaan et al., 2023). Our work complements these efforts by shifting the focus from eliciting reasoning to understanding the emergent mechanism of how this capability develops internally during training.
Complex Network Analysis of Neural Models. Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Estrada & Hatano, 2008). With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019; Ferrando et al., 2023) to better understand information flow. More recently, studies have explored how network structures emerge during supervised learning, identifying the formation of essential subnetworks (Li et al., 2023). A recent study by (Minegishi et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations. In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
@@ Appendix C The Reasoning Graph: From Empirical Extraction to a Unified Concept Web
Figure 7: Illustration of the LLM reasoning process. A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy ’forking tokens’ that represent decision points. The LLM reasoning process can be deconstructed into stable “concepts” (e.g., step 3) linked by high-entropy decision points. The probabilities at these junctures (e.g., “next” vs. “tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens.
@@
Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs. Their process involves generating a large corpus of reasoning chains, segmenting them into individual steps, and then computing a vector representation for each step by averaging the hidden states of its tokens at a specific Transformer layer. Finally, a clustering algorithm (e.g., K-means) is used on the full set of step vectors, with the resulting cluster centroids being defined as the graph’s nodes. By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance.
@@
While analyzing these static, task-specific graphs is insightful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks? It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web. This web represents the entirety of the model’s abstract knowledge, and the ultimate goal is to understand its structure.
@@
Figure 8: From Intractable LLM Dynamics to a Minimal CoNet Model. The reasoning process in an LLM (left) can be viewed as a traversal on a high-dimensional latent graph, where path probabilities are determined by the complex, autoregressive Transformer. Directly analyzing this graph’s evolution is intractable. CoNet (right) provides a minimal abstraction by replacing the intractable latent graph with a fixed K-regular random graph and simplifying the generation policy to a learnable Markov chain. This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation.
@@
However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles. The difficulties are multi-fold. First is a problem of scale: the true concept web is of astronomical size, and any extracted graph is an infinitesimal sample. More fundamentally is the ambiguity of the concept mapping: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one. Finally, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training. These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.

### v2

@@ 3 The Double-Edged Sword of Sparsity: from catastrophic Forgetting to Simulated Annealing scheme
From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability. It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with supervised learning (Li et al., 2024). Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I). As expected, the intervention triggered a sharp drop in overall performance in both CoNet and the 1.5B LLM (Fig. 44(a), 4(b)). Having reproduced this known macroscopic phenomenon, we leveraged CoNet’s transparency to inspect the microscopic aftermath.
The result is a direct validation of our structural explanation. The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web. Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, precisely breaking the single connection points for vast downstream subgraphs. Fig. 44(c)4(d) provides a stark visualization of this: a unified web is instantly fragmented, offering a clear, structural cause for the catastrophic performance drop.
@@ 4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
Although the overall reward curve in RLVR training looks smooth, our sparse-web theory shows that much more is happening underneath. As shown in the individual trajectories in Fig. 5, the network’s sparse structure creates two competing effects: (i) a forgetting process caused by frustration, which is strongest at the start of the slow-learning phase, and (ii) an opposite, phase-transition-like learning process that gradually builds and integrates knowledge during slow learning.
@@
Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality. (i) Frustration-Induced Forgetting: At the onset of slow learning [see inset in (a)], the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed. (ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green). The smoother gradients in the LLM (vs. the clean steps in CoNet) can be attributed to finite-size effects, as explained in the main text.
@@ 4.1 Frustration-Induced Forgetting: The Price of Integration
The first dynamic is driven by frustration-induced forgetting: a competitive process where the advancement of some skills forces the deterministic regression of others as the model integrates its “skill islands”. This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig. 5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025a).
This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where “skill islands” compete for limited connections under the $\left<k\right>\approx 2$ constraint. Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig. 6(b), blue). This subsequent degradation in performance is consistent with the phenomenon of policy collapse, which has been recently reported in several studies to occur after extended RLVR, leading to a loss of exploratory power (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
@@ 4.2 Phase-Transition-Like Learning: The Engine of Growth
The opposing dynamic, phase-transition-like learning, is an important engine of new knowledge acquisition (Power et al., 2022; Lin et al., 2025). The evidence for this is clearly shown in CoNet’s sharp, step-like accuracy jumps (Fig. 5(a)), and the steep accelerated learning gradients observed in the 1.5B LLM (Fig. 5(b), more details in Appendix G). This phenomenon is a multi-task extension of the “Learning at Criticality” (LaC) mechanism, first identified in a single-task context (Cai et al., 2025). The LaC framework models this process as a continuous phase transition (Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one. Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions. Beyond this transition, the policy collapses onto fixed optimal paths; the associated generalization drop is possibly linked to overfitting within the RLVR process (Dong et al., 2025; Gui & Cheng, 2025; Shypula et al., 2025; Zheng et al., 2025). This context explains the smoother gradient in the LLM. Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of finite-size effects (Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting. Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition. Shielded from interference, a new skill can be efficiently integrated via a clean phase transition. This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025). We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
5 Annealed-RLVR: An Algorithm Born from the Sparse-Web Theory
(a) Training Sample (512 Problems)
@@ 5 Annealed-RLVR: An Algorithm Born from the Sparse-Web Theory
(c) AIME 2024/2025
Figure 6: Annealed-RLVR Outperforms Standard RLVR. This figure compares the best@k accuracy curves of our Annealed-RLVR ("Annealed") with the standard RLVR baseline ("RLVR"). The evaluation is performed on (a) an in-distribution set (Randomly Selected 512 Training Problems) and two out-of-distribution (OOD) sets: (b) the Minerva dataset and (c) the AIME 2024/2025 datasets. The results show that Annealed-RLVR consistently achieves superior performance across all in-distribution and OOD benchmarks, confirming that the intervention fosters improved generalization.
@@
The algorithm operationalizes this concept with a precisely timed intervention. This intervention is triggered by monitoring the macroscopic signatures of the maximally frustrated state, which, as shown in Fig. 1, correspond to the crossover regime characterized by the “knee” of the reward curve (when accuracy gains plateau) and the bottom of the V-shaped response length curve. At this juncture, a brief SFT “heating” phase begins. The timing is critical: because the “skill islands” have not yet solidified into a single web, the system is uniquely robust to a targeted shock that would otherwise cause catastrophic forgetting—a theoretical claim confirmed by our experiments (Fig. 4). The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes. We implement this by surgically applying SFT only to problems with low accuracy (e.g., <0.1<0.1) for which correct solutions exist. Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
We validated this approach on the DeepSeek-R1-Distill-Qwen-1.5B model, trained on the DeepScaleR-Preview-Dataset . Our method, Annealed-RLVR, was compared against the standard RLVR baseline where the policy is optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024; Liu et al., 2025; Yu et al., 2025). The results (Fig. 6) confirm our theory. Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution evaluation set (512 randomly selected training problems) and on out-of-distribution (OOD) benchmarks, including the Minerva and AIME 2024/2025 dataset. The histograms (see Fig. 13) reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration. The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery. This results in a systematic reduction of unsolved problems and a corresponding population shift towards solved ones, confirming that our intervention guides the model past the inherent bottlenecks of standard RLVR to achieve superior reasoning. More details are given in Appendix H and I.
@@ 6 Conclusion and Outlook
A natural consequence of this sparse structure is a fundamental RLVR mechanism for forgetting and learning within the microscopic training process. We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently integrated at the web’s expanding frontier , yet this very per-skill collapse onto a single high-reward path also accumulates over time into a macroscopic policy collapse and a global loss of solution diversity when training proceeds too far beyond criticality.
@@ Appendix B Extended Related Work
System-2 Reasoning in LLMs. The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior. While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting deliberate System-2 reasoning. Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps. Other works aim to build this deliberative capacity into the model’s architecture or training process through self-correction and reflection (Shinn et al., 2023; Madaan et al., 2023). Our work complements these efforts by shifting the focus from eliciting reasoning to understanding the emergent mechanism of how this capability develops internally during training.
Complex Network Analysis of Neural Models. Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Simard et al., 2005; Estrada & Hatano, 2008). With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019) to better understand information flow (Abnar & Zuidema, 2020; Chefer et al., 2021). More recently, studies have explored the topological properties of neural networks, identifying optimal graph structures (You et al., 2020) and the formation of essential sparse subnetworks (Frankle & Carbin, 2019; Filan et al., 2021). Complementing these static analyses, recent work has modeled the training process itself as the evolution of complex networks (Danovski et al., 2024; Rocchetti, 2025). A recent study by (Wang et al., 2024; Minegishi et al., 2025; Matsutani et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations. In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
@@ Appendix C The Reasoning Graph: From Empirical Extraction to a Unified Concept Web
Figure 7: Illustration of the LLM reasoning process. A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy ’forking tokens’ that represent decision points. The LLM reasoning process can be deconstructed into stable “concepts” (e.g., step 3) linked by high-entropy decision points. The probabilities at these junctures (e.g., “next” vs. “tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens. This figure is adapted from Ref. Cai et al. (2025).
@@
Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs. Their process involves generating a large corpus of reasoning chains, segmenting them into individual steps, and then computing a vector representation for each step by averaging the hidden states of its tokens at a specific Transformer layer. Finally, a clustering algorithm (e.g., K-means) is used on the full set of step vectors, with the resulting cluster centroids being defined as the graph’s nodes. By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance. However, the cluster centroids derived from K-means offer limited interpretability as stable reasoning concepts, making their correspondence to the model’s internal processes unclear. Moreover, the induced graph’s structure is unstable, varying significantly with the choice of model, dataset, segmentation, and representation layer. Hence, this graph cannot track evolving representational dynamics during RLVR training process. How to construct a principled model of an LLM’s internal reasoning process remains a formidable open problem.
@@
While analyzing these static, task-specific graphs is useful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks? It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web. This web represents the entirety of the model’s abstract knowledge, and the ultimate goal is to understand its structure.
@@
Figure 8: From Intractable LLM Dynamics to a Minimal CoNet Model. The reasoning process in an LLM (left) can be viewed as a traversal on a high-dimensional latent graph, where path probabilities are determined by the complex, autoregressive Transformer. Directly analyzing this graph’s evolution is intractable. CoNet (right) provides a minimal abstraction by replacing the intractable latent graph with a fixed K-regular random graph and simplifying the generation policy to a learnable Markov chain. This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation. This figure is adapted from Ref. Cai et al. (2025).
@@
However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles. The difficulties are multifold. First, there is an issue of scale: the true concept web is of astronomical size, and any extracted graph is a small sample. Second, the mapping itself is ambiguous: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one. Third, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training. These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.